OpenStudy (anonymous):

the quotion of 3 times negitive 3 and the sum of negitive 3 and 2

OpenStudy (causticsyndicalist):

\[\left( 3 \times -3 \right) \div \left( -3 + 2 \right) = 9\] I hope I did it right.